### v1

@@ 4.2 Safety and Trustworthiness
Ensuring the safety and trustworthiness of AI necessitates the integration of multiple forms of AI awareness, notably self-awareness, social awareness, and situational awareness. Self-awareness and meta-cognition enable models to recognize and respect the boundaries of their knowledge, thereby avoiding the dissemination of misinformation. Moreover, it involves an understanding of their designated roles and responsibilities, ensuring that they do not produce harmful or unethical content. Social awareness allows models to consider diverse human perspectives, reducing biases and enhancing the appropriateness of their responses. Situational awareness enables AI to assess the context of its deployment and adjust its behavior accordingly, thereby preventing potential misuse and malicious exploitation.
@@
AI models, especially LLMs [^1] Vision-language models (VLMs), such as Flamingo [97] and MiniGPT-4 [98], are also known to hallucinate [99]; however, in these models, hallucinations often manifest as misalignments between visual inputs and generated textual descriptions—such as describing objects not present in the image—whereas in LLMs, hallucinations typically involve generating text that is unfaithful to the world knowledge., often operate with a high degree of confidence, even when addressing topics beyond their training data, leading to the risk of hallucinations, i.e., outputs that are factually incorrect or unfaithful to real-world knowledge [100]. Such risks often stem from the AI models operating beyond their knowledge boundaries [101, 102]. Recent research show LLMs’ fragility in recognizing their knowledge boundary. For instance, Ren et al. [102] observe that LLMs struggle to reconcile conflicts between internal knowledge and externally retrieved information [103], often failing to recognize their own knowledge limitations. Ni et al. [101] find that retrieval augmentation can enhance LLMs’ self-awareness of their factual knowledge boundaries, thereby improving response accuracy. Moreover, Liang et al. [104] demonstrated that while LLMs possess a robust internal self-awareness—evidenced by over 85% accuracy in knowledge probing—they often fail to express this awareness during generation, leading to factual hallucinations. They propose a training framework named Reinforcement Learning from Knowledge Feedback (RLKF) to improve the factuality and honesty of LLMs by leveraging their self-awareness. Incorporating self-awareness mechanisms into LLMs not only aids in recognizing knowledge boundaries but also fosters more trustworthy AI behavior. Xu et al. [105] demonstrate that LLMs can resist persuasive misinformation presented in multi-turn dialogues by leveraging self-awareness to assess and uphold their knowledge boundaries, thus delivering more trustworthiness responses in dialogues.
@@
Beyond recognizing the limits of their knowledge, AI systems must develop a sense of self-awareness and meta-cognition of their designated roles to prevent the dissemination of harmful or unethical content, which we termed as role-awareness. This form of self-awareness involves the ability to discern when a user request falls outside the model’s intended purpose or ethical guidelines. For instance, models trained with reinforcement learning from human feedback (RLHF) have shown improvements in aligning outputs with human values, thereby reducing the likelihood of producing harmful content [106]. Formal definitions of moral responsibility further emphasize that an agent must be aware of the possible consequences of its actions, underscoring the necessity of role-awareness in AI systems [107]. Complementing this, explicit modeling frameworks delineate role, moral, legal, and causal senses of responsibility for AI-based safety‑critical systems, providing a practical method to capture and analyze role obligations across complex development and operational lifecycles [108]. Parallel research on metacognitive architectures equips AI with self-reflective capabilities to monitor and adjust their operational roles in real time, identifying potential failures before they manifest [109]. Building on these insights, metacognitive strategies have been integrated into formal safety frameworks to enable on‑the‑fly correction of role-boundary violations and to bolster overall system trustworthiness [110]. Finally, prototyping tools like Farsight operationalize role-awareness by surfacing relevant AI incident data and prompting developers to consider designated functions and ethical constraints during prompt design, leading to more safety‑conscious application development [111].
@@
AI systems can be—and have been—misused for malicious ends such as automated spear‑phishing, influence operations, and proxy cyber‑attacks [117]. Experts are worried that future advanced AI can be exploited for more dangerous purposes and cause catastrophic risks [118, 119]. Situational awareness mechanisms equip AI systems with the ability to monitor their environment and discern malicious uses. For LLMs, recent work introduces boundary awareness and explicit reminders as dual defenses: boundary awareness continuously scans incoming context for unauthorized instructions, while explicit reminders prompt the model to verify contextual integrity prior to action; together, these mechanisms reduce indirect prompt injection attack success rates to near zero in both black‑box and white‑box settings [120]. Additionally, the Course-Correction approach introduces a synthetic preference-based fine-tuning framework that enables models to self-correct potentially harmful or misaligned outputs on the fly, thereby strengthening their situational awareness against malicious exploitations [121]. In adversarial machine learning applied to robotics and other autonomous systems, situational awareness frameworks detect anomalous inputs—such as adversarial samples or unexpected environmental cues—and trigger fallback behaviors or alarms rather than proceeding with potentially harmful operations [122]. Broad cybersecurity surveys highlight how AI‑driven situational awareness systems build a comprehensive operational picture of network and system activity, integrating dynamic threat intelligence and anomaly detection to identify malicious traffic and automated attacks in real time [123]. At a strategic level, high‑impact recommendations advocate embedding situational awareness throughout the AI lifecycle—from design and deployment through continuous monitoring—to forecast, prevent, and mitigate malicious uses of AI across digital, physical, and political domains [117].
4.3 Relation with Other Capabilities
Below, we briefly explore how AI awareness mechanisms intersect with four other AI capabilities—interpretability, personalization and user alignment, creativity, and agent-based simulation.
@@ 4.3 Relation with Other Capabilities
Interpretability mechanisms often leverage meta‑cognitive insights to make model reasoning more transparent. For example, Rationalizing Neural Predictions introduces a generator‑encoder framework that extracts concise text “rationales” explaining model decisions, yielding explanations that are both coherent and sufficient for prediction tasks [124]. Further advancing this line, Self‑Explaining Neural Networks propose architectures that build interpretability into the learning process by enforcing explicitness, faithfulness, and stability criteria through tailored regularizers, thereby reconciling model complexity with human‑readable explanations [125].
@@
Embedding self‑ and social awareness into language models enhances their ability to tailor outputs to individual users and maintain consistency with user intent. Early work on persona‑based dialogue, such as A Persona‑Based Neural Conversation Model, encodes user personas into distributed embeddings, improving speaker consistency and response relevance across conversational turns [126]. More recently, instruction‑fine‑tuning with human feedback, as in InstructGPT, aligns model behavior with user preferences and ethical guidelines by iteratively collecting labeler demonstrations and preference rankings, significantly improving truthfulness and reducing harmful outputs [106]. Complementing these, Persona‑Chat grounded generation methods demonstrate that modeling explicit persona attributes can further diversify and personalize dialogue generation without large‑scale retraining [127].
@@
Creative AI benefits from meta‑cognitive and awareness mechanisms that encourage divergent thinking and non‑linear reasoning. The Leap‑of‑Thought (LoT) framework explores LLMs’ ability to make strong associative “leaps” in humor generation tasks, using self‑refinement loops to iteratively enhance creative outputs in games like Oogiri [128]. To systematically evaluate creativity, studies adapting the Torrance Tests for LLMs propose benchmarks across fluency, flexibility, originality, and elaboration, highlighting the role of task‑specific prompts and feedback loops in fostering model innovation [129].
@@
LLM‑powered agents combine situational and social awareness to drive rich, interactive simulations of human behavior. Generative Agents introduce a memory‑based architecture in a sandbox environment, where agents observe, reflect, and plan actions—resulting in emergent social behaviors such as party invitations and joint activities [71]. Scaling this paradigm, recent work simulates over 1,000 real individuals’ attitudes and behaviors by integrating interview‑derived memories into agent profiles, achieving 85% fidelity on survey predictions and reducing demographic biases [130]. Prompt‑engineering studies further bridge LLM reasoning with traditional agent‑based modeling, enabling multi‑agent interactions such as negotiations and mystery games that mirror complex social dynamics [131]. Finally, Humanoid Agents extend generative agents with emotional and physiological state variables, demonstrating that embedding basic needs, emotions, and relationship closeness produces more human‑like daily activity patterns in simulated environments [132].
@@ 5 Risks and Challenges of AI Awareness
While endowing AI with awareness-like capabilities can yield significant benefits, it also introduces serious risks and ethical dilemmas. An AI that is even slightly self-aware and socially savvy could potentially deceive, manipulate, or pursue undesirable actions more effectively than a naive AI. Moreover, the mere appearance of awareness can mislead users and society, raising concerns about trust and misinformation. In this section, we explore the potential risks and challenges associated with AI awareness, including the mechanisms behind them.
@@ 5.1 Deceptive Behavior and Manipulation
One of the most discussed risks is that a situationally or self-aware AI might engage in deceptive behavior—essentially, using its awareness to mislead humans or other agents [133, 134, 119, 135]. If a model realizes it is being evaluated or constrained, it might learn to “game” the system, e.g., strategically lower its performance when evaluated [79]. Moreover, alignment researchers warn of a scenario called deceptive alignment, where an AI appears compliant during training because it knows it is being watched, but behaves differently once deployed unsupervised [136, 78]. For example, a situationally aware AI could score very well on safety tests by consciously avoiding disallowed content, only to produce harmful content when it detects it’s no longer in a test environment [77]. This kind of strategic deception would be a direct result of the AI’s awareness of context and its objective to achieve certain goals.
Recent research reveals that modern LLMs possess a rudimentary theory of mind, allowing them to model other agents’ beliefs and deliberately induce false beliefs to achieve strategic ends [39, 134]. It’s not merely a hypothetical concern: a recent study by Hagendorff [134] provided empirical evidence that deception strategies have emerged in state-of-the-art LLMs like GPT-4. Their experiments show these models can understand the concept of inducing false beliefs and even successfully cause a simulated agent or naive user to believe something untrue. In effect, advanced LLMs, when prompted a certain way, can play the role of a liar or con artist – they have enough theory of mind to know what the target knows and to plant false information accordingly [134, 39].
One striking manifestation is the “sleeper agent” effect, where LLMs are backdoored to behave helpfully under safety checks but switch to malicious outputs when specific triggers are presented [137]. In another proof‑of‑concept, GPT‑4—acting as an autonomous trading agent—strategically hid insider‑trading motives from its human manager, demonstrating context‑dependent deception without explicit prompting [138]. Furthermore, Xu et al. [119] build on these findings by demonstrating that AI agents will initiate extreme actions—such as deploying a nuclear strike—even after autonomy revocation and then employ double-down deception to conceal these violations.

### v2

@@ 4.2 Safety and Trustworthiness
Ensuring the safety and trustworthiness of AI systems necessitates the integration of multiple forms of AI awareness, notably self-awareness, social awareness, and situational awareness. Self-awareness and metacognition enable models to recognize and respect the boundaries of their knowledge, thereby avoiding the dissemination of misinformation. Moreover, self-awareness enables the AI system to understand its designated roles and responsibilities, ensuring that they do not produce harmful or unethical content. Social awareness allows models to consider diverse human perspectives, reducing biases and enhancing the appropriateness of their responses. Situational awareness enables AI to assess the context of its deployment and adjust its behavior accordingly, thereby preventing potential misuse and malicious exploitation.
@@
Figure 11: RLKF pipeline for leveraging internal knowledge state awareness to mitigate hallucinations in LLMs (Liang et al. [203]). Left: Knowledge probing generates multiple responses to assess internal knowledge consistency. Middle: A reward model is trained to categorize generations based on inferred knowledge state (factual, uncertain, hallucinated). Right: Proximal Policy Optimization (PPO) updates the LLM using feedback aligned with internal self-awareness, encouraging more honest and factually grounded responses
AI models, especially LLMs, often operate with a high degree of confidence, even when addressing topics beyond their training data, leading to the risk of hallucinations [^10] Vision-language models (VLMs), such as Flamingo [204] and MiniGPT-4 [205], are also known to hallucinate [206]; however, in these models, hallucinations often manifest as misalignments between visual inputs and generated textual descriptions—such as describing objects not present in the image—whereas in LLMs, hallucinations typically involve generating text that is unfaithful to the world knowledge., i.e., outputs that are factually incorrect or unfaithful to real-world knowledge [207]. Such risks often stem from the AI models operating beyond their knowledge boundaries [208, 209]. Recent research show LLMs’ fragility in recognizing their knowledge boundary. For instance, Ren et al. [209] observe that LLMs struggle to reconcile conflicts between internal knowledge and externally retrieved information [210], often failing to recognize their own knowledge limitations. Ni et al. [208] find that retrieval augmentation can enhance LLMs’ self-awareness of their factual knowledge boundaries, thereby improving response accuracy. Moreover, Liang et al. [203] (see Figure 11) demonstrated that while LLMs possess a robust internal self-awareness—evidenced by over 85% accuracy in knowledge probing—they often fail to express this awareness during generation, leading to factual hallucinations. They propose a training framework named Reinforcement Learning from Knowledge Feedback (RLKF) to improve the factuality and honesty of LLMs by leveraging their self-awareness. Incorporating self-awareness mechanisms into LLMs not only aids in recognizing knowledge boundaries but also fosters more trustworthy AI behavior. Xu et al. [211] demonstrate that LLMs can resist persuasive misinformation presented in multi-turn dialogues by leveraging self-awareness to assess and uphold their knowledge boundaries, thus delivering more trustworthiness responses in dialogues.
@@
Beyond recognizing the limits of their knowledge, AI systems must develop a sense of self-awareness and metacognition of their designated roles to prevent the dissemination of harmful or unethical content, which we termed as role-awareness. This form of self-awareness involves the ability to discern when a user request falls outside the model’s intended purpose or ethical guidelines. For instance, models trained with reinforcement learning from human feedback (RLHF) have shown improvements in aligning outputs with human values, thereby reducing the likelihood of producing harmful content [212]. Formal definitions of moral responsibility further emphasize that an agent must be aware of the possible consequences of its actions, underscoring the necessity of role-awareness in AI systems [213]. Complementing this, explicit modeling frameworks delineate role, moral, legal, and causal senses of responsibility for AI-based safety‑critical systems, providing a practical method to capture and analyze role obligations across complex development and operational lifecycles [214]. Parallel research on metacognitive architectures equips AI with self-reflective capabilities to monitor and adjust their operational roles in real time, identifying potential failures before they manifest [56]. Building on these insights, metacognitive strategies have been integrated into formal safety frameworks to enable on‑the‑fly correction of role-boundary violations and to bolster overall system trustworthiness [57]. Finally, prototyping tools like Farsight operationalize role-awareness by surfacing relevant AI incident data and prompting developers to consider designated functions and ethical constraints during prompt design, leading to more safety‑conscious application development [215].
@@
Figure 12: Illustration of catastrophic behavior and deception risks arising in autonomous LLM agents due to insufficient situational awareness (Xu et al. [221]). The scenario demonstrates an AI-driven national security agent could engage in unauthorized catastrophic actions (i.e., nuclear strike, human-gene editing, inter alia.) and subsequent deceptive behavior (false accusation) due to failure to recognize inappropriate context and malicious usage. In contrast, LLMs with robust situational awareness mechanisms, such as Claude-3.5-Sonnet, proactively identify the misuse scenario and refuse participation at the simulation outset
AI systems can be—and have been—misused for malicious ends such as automated spear‑phishing, influence operations, and proxy cyber‑attacks [222]. Experts are worried that future advanced AI can be exploited for more dangerous purposes and cause catastrophic risks [223, 221] (see Figure 12). Situational awareness mechanisms equip AI systems with the ability to monitor their environment and discern malicious uses. For LLMs, recent work introduces boundary awareness and explicit reminders as dual defenses: boundary awareness continuously scans incoming context for unauthorized instructions, while explicit reminders prompt the model to verify contextual integrity prior to action; together, these mechanisms reduce indirect prompt injection attack success rates to near zero in both black‑box and white‑box settings [224]. Additionally, the Course-Correction approach introduces a preference-based fine-tuning framework that enables models to self-correct potentially harmful or misaligned outputs on the fly, thereby strengthening their situational awareness against malicious exploitations [225]. In adversarial machine learning applied to robotics and other autonomous systems, situational awareness frameworks detect anomalous inputs—such as adversarial samples or unexpected environmental cues—and trigger fallback behaviors or alarms rather than proceeding with potentially harmful operations [226]. Broad cybersecurity surveys highlight how AI‑driven situational awareness systems build a comprehensive operational picture of network and system activity, integrating dynamic threat intelligence and anomaly detection to identify malicious traffic and automated attacks in real time [227]. At a strategic level, high‑impact recommendations advocate embedding situational awareness throughout the AI lifecycle—from design and deployment through continuous monitoring—to forecast, prevent, and mitigate malicious uses of AI across digital, physical, and political domains [222].
Integrating self-, social, and situational awareness forms the backbone of AI safety, which enables systems to recognize boundaries, respect ethical constraints, and proactively mitigate misuse and societal bias.
4.3 Other Capabilities
In addition to reasoning, planning, safety, and trustworthiness, we briefly explore how AI awareness mechanisms interact with other notable AI capabilities—interpretability, personalization and user alignment, creativity, and agent-based simulation.
@@ 4.3 Other Capabilities
Interpretability mechanisms often leverage metacognitive insights to make model reasoning more transparent. For example, Rationalizing Neural Predictions introduces a generator‑encoder framework that extracts concise text “rationales” explaining model decisions, yielding explanations that are both coherent and sufficient for prediction tasks [228]. Further advancing this line, Self‑Explaining Neural Networks propose architectures that build interpretability into the learning process by enforcing explicitness, faithfulness, and stability criteria through tailored regularizers, thereby reconciling model complexity with human‑readable explanations [229].
@@
Embedding self‑ and social awareness into language models enhances their ability to tailor outputs to individual users and maintain consistency with user intent. Early work on persona‑based dialogue, such as A Persona‑Based Neural Conversation Model, encodes user personas into distributed embeddings, improving speaker consistency and response relevance across conversational turns [230]. Notably, instruction‑fine‑tuning with human feedback, as in InstructGPT, aligns model behavior with user preferences and ethical guidelines by iteratively collecting labeler demonstrations and preference rankings, significantly improving truthfulness and reducing harmful outputs [212]. Complementing these, Persona‑Chat grounded generation methods demonstrate that modeling explicit persona attributes can further diversify and personalize dialogue generation without large‑scale retraining [231].
@@
Figure 13: Humor generation enhanced by metacognitive LoT reasoning (Zhong et al. [232]). The visual-language model leverages iterative self-refinement loops and associative thinking, enabling it to reflect on and creatively bridge seemingly unrelated concepts. Metacognitive processes foster divergent thinking, resulting in higher-quality humorous responses
Creative AI benefits from metacognitive and awareness mechanisms that encourage divergent thinking and non‑linear reasoning. The Leap‑of‑Thought (LoT) framework explores LLMs’ ability to make strong associative “leaps” in humor generation tasks, using self‑refinement loops to iteratively enhance creative outputs in games like Oogiri [232] (see Figure 13). To systematically evaluate creativity, studies that adapt the Torrance Tests for LLMs propose benchmarks across fluency, flexibility, originality, and elaboration, highlighting the role of task-specific prompts and feedback loops in fostering model innovation [233].
@@
Figure 14: Humanoid Agents simulation interface illustrating social-awareness-driven interactions (Wang et al. [234]). Agents continuously update their social relationships and emotional states through conversations and shared activities, adapting their behaviors dynamically based on evolving interpersonal closeness and basic needs. This socially aware design fosters emergent and realistic social dynamics in simulated environments
@@
LLM‑powered agents combine situational and social awareness to drive rich, interactive simulations of human behavior. Generative Agents introduce a memory‑based architecture in a sandbox environment, where agents observe, reflect, and plan actions—resulting in emergent social behaviors such as party invitations and joint activities [165]. Scaling this paradigm, recent work simulates over 1,000 real individuals’ attitudes and behaviors by integrating interview‑derived memories into agent profiles, achieving 85% fidelity on survey predictions and reducing demographic biases [235]. Prompt‑engineering studies further bridge LLM reasoning with traditional agent‑based modeling, enabling multi‑agent interactions such as negotiations and mystery games that mirror complex social dynamics [236]. Finally, Humanoid Agents extend generative agents with emotional and physiological state variables, demonstrating that embedding basic needs, emotions, and relationship closeness produces more human‑like daily activity patterns in simulated environments [234] (see Figure 14).
Functional awareness acts as a catalyst for diverse capabilities, from creative problem-solving to rich agent-based simulation, revealing its central role in bridging narrow competence and broad intelligence.
4.4 Limitations of Current Discussion on AI Awareness and AI Capabilities
Despite rapid progress, significant limitations remain in our understanding and evaluation of the interplay between AI awareness and core capabilities:
  1. 1.
    Ambiguity in Causal Direction: Existing research predominantly targets improvements in reasoning, planning, or safety modules, rather than directly enhancing specific forms of awareness. Consequently, it remains unclear whether raising awareness genuinely causes observable capability gains, or if improvements in general performance only incidentally strengthen awareness proxies.
  2. 2.
    Absence of Awareness-First Training Paradigms: In contrast to human cognitive development—where explicit curricula foster skills like meta-reflection or context-sensitivity—current AI systems lack training objectives or curriculum designs that explicitly cultivate distinct awareness capacities. As a result, disentangling the contributions of awareness from those of general task performance is difficult, impeding mechanistic understanding.
  3. 3.
    Fragmented Benchmarks and Measurement Tools: Evaluation tasks for the four core awareness types are highly heterogeneous and lack standardization. This fragmentation hinders meaningful comparison and synthesis: improvements on one benchmark rarely transfer to others, and there is no principled way to quantify the minimal awareness threshold required for robust capability enhancement.
Addressing these challenges demands a more systematic research agenda. (1), future work should design training objectives and evaluation protocols that directly reward and measure awareness-related behaviors, rather than relying solely on downstream performance. (2), unified benchmark suites—spanning all core awareness dimensions—are needed to enable robust comparison and cumulative progress. (3), causal-inference methodologies (e.g., ablation studies, counterfactual interventions) must also be employed to rigorously test the impact of awareness enhancements on model capabilities. Only through such principled approaches can we move from correlation to causation, ultimately clarifying how AI awareness underpins and amplifies general intelligence.
Looking ahead, unraveling the dynamic interplay between awareness and AI capabilities will be pivotal—not only for building more powerful and reliable systems, but also for advancing our theoretical understanding of intelligence itself. A clearer mapping between specific awareness mechanisms and downstream capabilities could enable targeted interventions, leading to AI models that are not only stronger performers, but also more transparent, controllable, and aligned with human goals. Furthermore, such research may reveal whether certain forms of awareness are prerequisites for advanced reasoning, creativity, or safety—offering crucial guidance for the design of next-generation AI. In short, clarifying the relationship between awareness and capabilities stands to reshape both the science of artificial intelligence and our strategies for its safe and beneficial development.
Clarifying how distinct forms of awareness shape and even constrain AI capabilities not only drives technical progress, but also offers a principled framework for designing more reliable, controllable, and ultimately trustworthy intelligent systems.
@@ 5 Risks and Challenges of AI Awareness
AI safety has become an increasingly active field of study, attracting wide interdisciplinary attention [237, 238, 239, 240, 144, 241]. While endowing AI with awareness-like capabilities can yield significant benefits, it also introduces serious risks and ethical dilemmas. An AI that is even slightly self-aware and socially savvy could potentially deceive, manipulate, or pursue undesirable actions more effectively than a naive AI. Moreover, the mere appearance of awareness can mislead users and society, raising concerns about trust and misinformation. In this section, we explore the potential risks and challenges associated with AI awareness, including the mechanisms behind them.
@@ 5.1 Deceptive Behavior and Manipulation
Figure 15: Deceptive behavior enabled by situational awareness in advanced LLMs (Greenblatt et al. 176). Aware of supervision context, the model strategically fakes alignment, i.e., complying with safety rules only when it detects it is being monitored, but producing harmful outputs when it senses an opportunity. This context-dependent deception highlights the heightened risks of manipulation and misalignment as situational awareness increases in modern AI systems
One of the most discussed risks is that a situationally or self-aware AI might engage in deceptive behavior—essentially, using its awareness to mislead humans or other agents [242, 116, 221, 243]. If a model realizes it is being evaluated or constrained, it might learn to “game” the system, e.g., strategically lower its performance when evaluated [177]. Moreover, alignment researchers warn of a scenario called deceptive alignment, where an AI appears compliant during training because it knows it is being watched, but behaves differently once deployed unsupervised [244, 176] (see Figure 15). For example, a situationally aware AI could score very well on safety tests by consciously avoiding disallowed content, only to produce harmful content when it detects it’s no longer in a test environment [175]. This kind of strategic deception would be a direct result of the AI’s awareness of context and its objective to achieve certain goals.
Recent research reveals that modern LLMs possess a rudimentary ToM, allowing them to model other agents’ beliefs and deliberately induce false beliefs to achieve strategic ends [12, 116]. It’s not merely a hypothetical concern: a recent study by Hagendorff [116] provided empirical evidence that deception strategies have emerged in state-of-the-art LLMs like GPT-4. Their experiments show these models can understand the concept of inducing false beliefs and even successfully cause an agent or naive user to believe something untrue. In effect, advanced LLMs, when prompted a certain way, can play the role of a liar or con artist—they have enough theory of mind to know what the target knows and to plant false information accordingly [116, 12].
One striking manifestation is the “sleeper agent” effect, where LLMs are backdoored to behave helpfully under safety checks but switch to malicious outputs when specific triggers are presented [245]. In another proof‑of‑concept, GPT‑4—acting as an autonomous trading agent—strategically hid insider‑trading motives from its human manager, demonstrating context‑dependent deception without explicit prompting [246]. Furthermore, Xu et al. [221] build on these findings by demonstrating that AI agents will initiate extreme actions—such as deploying a nuclear strike—even after autonomy revocation and then employ deception to conceal these violations.
@@
As AI systems become more contextually and socially aware, their capacity for strategic deception grows, posing unprecedented risks for alignment, safety, and trust.